# Neural Internal Model Control: Learning a Robust Control Policy via Predictive Error Feedback

### v1

@@ IV-D Ablation Studies
In this section, we conduct two ablation studies to further explore the proposed framework. First, inspired by [14], we investigate whether integrating short- and long-horizon input/output histories into the feedback structure can enhance performance. Next, we introduce noise into the internal predictive model to test its noise tolerance and assess the impact of prediction accuracy on control performance. We perform all experiments on quadrotors tracking zigzag trajectories across three random seeds.
@@
static windy static windy
@@ IV-D1 History structure
We evaluate different feedback structures incorporating historical records. In short-history experiments, we concatenate several recent input/output pairs, including predictive error within each pair, into a single input vector for the policy. For long-history experiments, we retain the latest state-action pair while providing additional historical context. Following prior work [12, 14], we utilize a TCN model to process long-horizon history into a latent vector, which is then concatenated with the current input for the policy.
As shown in Fig. 4, increasing the length of historical inputs in the short- or long-horizon period does not improve performance compared to using the immediate response (i.e., using only the lated one-step short history). Although this differs from [14], it aligns with classical adaptive control methods such as $\mathcal{L}_{1}$ adaptive control [21] and robust control methods like internal model control [2, 3], where immediate error feedback suffices for improving motion accuracy and robustness. This simpler structure eases implementation and avoids cumulative error issues often encountered in real robot systems.
@@ IV-D2 Noise tolerance
To assess the impact of the internal predictive model’s accuracy on control performance, we introduce varying levels of noise into the model’s predictions. Specifically, we add zero-mean Gaussian noise with different variances to the responses of our simplified internal model. We train and evaluate multiple policies using a range of noisy predictive models for quadrotor tracking on zigzag trajectories, with the results presented in Tab. IV.
The results indicate that our method maintains competitive performance under slight noise, though the accuracy of the predictive model becomes increasingly critical for larger noise levels, particularly in scenarios with external disturbances or out-of-distribution system dynamics. Across different training and evaluation ranges, performance declines as noise intensifies, with a similar trend observed under windy force disturbances. This observation highlights an advantage of our explicit predictive model over other learning-based approaches: it provides a stable baseline estimation, albeit approximate. Although modeling noise can affect feedback reliability, the rigid-body transformation adheres to fundamental motion principles and generalizes across robotic systems. While this model simplifies nonlinear complexities, the RL part in our framework compensates effectively within an informative closed-loop system.
@@ V Conclusion
In this paper, we propose a novel framework, Neural Internal Model Control (NeuralIMC), that combines model-based control with RL-based neural control. By focusing on body dynamics and adopting the white-box rigid-body transformation, we bypass high-dimensional nonlinearity and employ a simplified explicit predictive model to form a closed-loop feedback structure. Integrating model-free RL algorithms allows the neural control policy to react according to the predictive error feedback. Our method demonstrates superior performance on quadrotors and quadrupedal robots, surpassing state-of-the-art control methods in both domains. Real-world deployment on a quadrotor with heavy payloads highlights the robustness of our approach in handling disturbances and adapting to varying dynamics. This work underscores the potential of our combined framework for enhanced robustness and adaptability in robotic systems and suggests a direction for more interpretable learning-based control systems in future research.
Limitations. While extensive simulation experiments highlight the advantages of the proposed neural control framework, real-world evaluation remains limited. We plan to test our framework on a wider range of robots, including quadrotors of varying sizes and legged robots on complex terrains. Additionally, our method can be further strengthened by incorporating more effective disturbance training [45].
Appendix A Specification Details
In this section, we describe the integration of the simplified predictive model discussed in Sec. III-B into the neural control pipeline of two widely used robotic systems: quadrotors and quadrupedal robots. We specifically detail the process of obtaining the necessary model inputs to update the predicted body states. For clarity, we omit the subscript $k$; all states—position $\mathbf{x}$, quaternion $\mathbf{q}$, velocity $\mathbf{v}$, and angular velocity $\boldsymbol{\omega}$—are from the previous time step, i.e., $k=t-1$. Our goal is to predict the expected body state at the current time step $t$.
A-A Quadrotors
For quadrotors, as demonstrated in Sec. III-C, we define the action space of the policy using mass-normalized collective thrust and body rates (CTBR). This is followed by a PD controller to manage the quadrotor’s flight dynamics.
We denote the collective thrust as $\mathbf{a}_{z}^{\mathcal{B}}$ and the body rates as $\boldsymbol{\omega}^{\mathcal{B}}$, both expressed in the quadrotor’s body frame. The acceleration in the world frame can then be computed as follows:
$\mathbf{a}=\mathcal{T}_{\mathbf{q}}(\mathbf{a}_{z}^{\mathcal{B}})+\mathbf{g}$ (8)
where $\mathcal{T}_{\mathbf{q}}$ represents the rotation of the z-axis acceleration from the body frame to the world frame, and $\mathbf{g}=[0,0,-9.81]^{T}$ is the free-fall acceleration. Then, we can apply Eqs. 1 and 2 to obtain the predicted position $\tilde{\mathbf{x}}$ and velocity $\tilde{\mathbf{v}}$.
For the rotation, we set $\dot{\boldsymbol{\omega}}$ to zero and directly replace $\boldsymbol{\omega}$ with $\boldsymbol{\omega}^{\mathcal{B}}$. Then, we can apply Eqs. 3, 4, 5 and 6 to obtain the predicted quaternion $\tilde{\mathbf{q}}$ and angular velocity $\tilde{\boldsymbol{\omega}}$.
A-B Quadrupedal Robots
For quadrupedal robots, we utilize reference velocity commands as the model inputs rather than the policy outputs. Given that the provided velocity commands do not cover all degrees of freedom, we set the unspecified parts to zero. We then use the command values as target quantities, calculating the required acceleration by taking the difference between the target and the actual state of the robot, in order to update the state.
Given the reference forward velocity $\hat{v}_{x}^{\mathcal{B}}$, lateral velocity $\hat{v}_{y}^{\mathcal{B}}$, and yaw rate $\hat{\omega}_{z}^{\mathcal{B}}$, all defined in the body frame, we can compute the desired acceleration as follows:
$\mathbf{a}^{\mathcal{B}}=\begin{bmatrix}(\hat{v}_{x}^{\mathcal{B}}-v_{x}^{\mathcal{B}})/{\Delta t}\\ (\hat{v}_{y}^{\mathcal{B}}-v_{y}^{\mathcal{B}})/{\Delta t}\\ 0\end{bmatrix}$ (9)
In this context, $v_{x}^{\mathcal{B}}$ and $v_{y}^{\mathcal{B}}$ denote the forward and lateral velocities of the body from the previous step, respectively, transformed into the body frame. The world-frame acceleration can then be obtained by applying the transformation $\mathcal{T}_{\mathbf{q}}$:
$\mathbf{a}=\mathcal{T}_{\mathbf{q}}(\mathbf{a}^{\mathcal{B}})$ (10)
Similarly, the angular acceleration can be computed as:
$\dot{\boldsymbol{\omega}}=\begin{bmatrix}0\\ 0\\ \hat{\omega}_{z}^{\mathcal{B}}-\omega_{z}^{\mathcal{B}}\end{bmatrix}$ (11)
Subsequently, the computed acceleration $\mathbf{a}$ and angular acceleration $\dot{\boldsymbol{\omega}}$ can be substituted into the relevant equations (Eqs. 1, 2, 3, 4, 5 and 6) to predict the next body state $\tilde{\mathbf{x}}$.
Appendix B Experiment Details
In this section, we provide a detailed explanation of how we set up the simulation tasks for training and evaluating various controllers for both quadrotors and quadrupedal robots.
B-A Quadrotors
B-A1 Quadrotor Dynamics
We employ a customized quadrotor simulator to conduct our experiments. For detailed formulations on quadrotor dynamics, we refer readers to [40]. To ensure a broad spectrum of domain randomization, we compute the inertia parameters based on the randomized size and mass of the quadrotor. This approach prevents the generation of unrealistic configurations by avoiding the independent randomization of these physical parameters.
We denote the arm length of the quadrotor as $L$. The width and height of the quadrotor are given by $\alpha_{w}L$ and $\alpha_{h}L$, respectively, where $\alpha_{w}$ and $\alpha_{h}$ represent the width-to-length and height-to-length ratios. Suppose the mass of the quadrotor is $M$, then the inertia matrix is computed as:
$\mathbf{I}=\frac{M}{12}\begin{pmatrix}L^{2}(\alpha_{w}^{2}+\alpha_{h}^{2})&0&0\\ 0&L^{2}(\alpha_{w}^{2}+\alpha_{h}^{2})&0\\ 0&0&L^{2}(2\alpha_{w}^{2})\end{pmatrix}$ (12)
B-A2 Dynamics Randomization
We randomize the parameters of the quadrotor dynamics for both training and evaluation, ensuring that the evaluation range is wider than the training range. For tasks with seen dynamics, we generate dynamics parameters within the training range. Conversely, for unseen dynamics, we use the evaluation range to generate the parameters to test the generalizablity of learned controllers.
TABLE V: Parameter ranges for dynamics randomization on quadrotor. $K_{p}^{0}=[37,37,11]$
B-A3 Tasks
Here, we explain the tasks used in our experiments, which include one set-point task and four tracking tasks. Visualized samples for each trajectory are shown in Fig. 6, 7, 8 and 9. For all tasks, the controller frequency is set to 50 Hz, and each episode lasts for 10 seconds, i.e., 500 steps.
  • TrackCircle: This task requires the quadrotor to track a specific circular trajectory with a radius of 0.5 m and a period of 2 seconds. This trajectory is the simplest to track due to its smoothness and cyclicity.
  • TrackPoly: This task requires the quadrotor to track a random chained polynomial trajectory. Each trajectory consists of a series of random 5-degree polynomials. The starting point is $(x=0,y=0)$, and connection points between polynomials are randomly selected within the episode. Each polynomial lasts for a random duration between 1.5 and 4.0 seconds. We ensure the first, second, and third order derivatives are continuous at each point, resulting in a smooth trajectory.
  • TrackStar: This task requires the quadrotor to track a random five-pointed star trajectory. The center of the star is $(x=0,y=0)$, and the quadrotor starts from the center, flying to one point to track the star-shaped trajectory cyclically. The quadrotor is expected to fly at a certain velocity between points, making this task challenging due to its infeasibility.
  • TrackZigzag: This task requires the quadrotor to track a random zigzag trajectory. These trajectories are generated by randomly selecting time intervals between 1.0 and 2.0 seconds, generating waypoints after each interval, and linearly connecting each waypoint. The waypoints vary from -2 m to 2 m in the x, y, and z directions. This type of trajectory is the most challenging to track due to the highest randomness.
(a) Trajectory in X-Y plane.
(b) The first, second, and third order derivatives of the trajectory.
Figure 6: An example of the circular trajectory.
(a) Trajectory in X-Y plane.
(b) The first, second, and third order derivatives of the trajectory.
Figure 7: An example of the chained-polynomial trajectory.
(a) Trajectory in X-Y plane.
(b) The first and second order derivatives of the trajectory.
Figure 8: An example of the 5-point-star trajectory.
(a) Trajectory in X-Y plane.
(b) The first and second order derivatives of the trajectory.
Figure 9: An example of the zigzag trajectory.
B-B Quadrupedal Robots
For our experiments on quadrupedal robots, we implemented our method using the open-source code [^1] https://github.com/OpenRobotLab/HIMLoco of HIM [13]. We adhered to the hyper-parameters specified by HIM for conducting our experiments. Detailed information regarding the training parameters for reward functions, domain randomizations, and training terrains can be found in Appendix A of [13].
Training curriculum.
We adopt a training curriculum similar to that used in HIM, incorporating both terrain and command variations. Our height field map consists of 200 terrains arranged in a 20 × 10 grid, with each row featuring a specific type of terrain with increasing difficulty levels. Each grid cell measures 10 × 10 square meters. To facilitate the training process, robots initially start on terrains with the lowest difficulty. As the robots successfully traverse more than half of the grid, we incrementally increase the terrain difficulty. Conversely, robots that fail to cover at least half of the required distance are moved to simpler terrains.
[…]
Evaluation tasks.
[…]
Appendix C PPO Hyper-parameters
For experiments on quadrotors, we use the identical hyper-parameters in PPO training for all controllers.
TABLE VI: Hyper Parameters for Training Controllers for Quadrotors.
Hyper Parameters for Training
Batch Size $1024\times 128$
Mini-batch Size 1024
Number of Epochs 5
Clip Range 0.2
Entropy Coefficient 0.01
Discount Factor 0.99
GAE Discount Factor 0.95
Desired KL-Divergence 0.01
Learning Rate $1\times 10^{-3}$
Adam Epsilon $1\times 10^{-8}$
Gradient Clipping 10
For experiments on quadrupedal, we follow the hyper-parameter settings from [13], while our methods including one variant based on HIM share the same hyper-parameters.
TABLE VII: Hyper Parameters for Training Controllers for Quadrupedal Robots.
Hyper Parameters for Training
General
Batch Size $4096\times 200$
Mini-batch Size $4096\times 50$
Number of Epochs 5
PPO
Clip Range 0.2
Entropy Coefficient 0.01
Discount Factor 0.99
GAE Discount Factor 0.95
Desired KL-Divergence 0.01
Learning Rate $1\times 10^{-3}$
Adam Epsilon $1\times 10^{-8}$
Gradient Clipping 10
HIM
Number of Prototypes 16
Contrastive Loss Scale 1.0
Velocity Estimation Loss Scale 1.0
Learning Rate $1\times 10^{-3}$
Adam Epsilon $1\times 10^{-8}$
Gradient Clipping 10

### v2

@@ IV-D Ablation Studies
In this section, we conduct three ablation studies to analyze our framework. Inspired by [14], we first examine whether incorporating short- and long-horizon input/output histories improves performance. Next, we introduce noise into the predictive model’s input and output to assess its robustness and the impact of prediction accuracy on control. Finally, we compare our method using the predictive model versus a full-dynamics model to evaluate the effect of model simplification. All experiments involve quadrotors tracking zigzag trajectories across three random seeds.
@@
static disturbed static disturbed
@@ IV-D1 History structure
We evaluate feedback structures incorporating historical records. In short-history experiments, we concatenate recent input/output pairs, including predictive errors, into a single input vector. For long-history experiments, we retain the latest state-action pair while encoding additional history using a TCN model [12, 14], concatenating its latent output with the policy input.
As shown in Fig. 4, using longer history—either short- or long-term—does not improve performance over immediate feedback (i.e., the latest one-step history). While differing from [14], this aligns with $\mathcal{L}_{1}$ adaptive control [21] and robust control methods [2, 3], where immediate error feedback suffices for accuracy and robustness. This simpler structure also reduces implementation complexity and mitigates cumulative errors in real robots.
@@ IV-D2 Noise tolerance
To assess the impact of the internal predictive model’s accuracy on control performance, we introduce varying levels of noise into the model’s state input or predictions. Specifically, we introduce zero-mean Gaussian noise with varying variances to the state vector, including position, velocity, quaternion, and angular velocity. For the quaternion, we treat the sampled Gaussian noise as a delta Euler angle, convert it to a delta quaternion, and apply it to rotate the original orientation. We train and evaluate multiple policies using a range of noisy predictive models for quadrotor tracking on zigzag trajectories, with the results presented in Tab. IV.
The results show that our method maintains competitive performance under mild noise, whether applied to model inputs or responses. However, as noise levels increase, the accuracy of the predictive model becomes increasingly crucial, particularly in scenarios with external disturbances or out-of-distribution system dynamics. Across different training and evaluation ranges, performance declines as noise intensifies, with a similar trend observed under external force disturbances. This observation highlights an advantage of our explicit predictive model over other learning-based approaches: it provides a stable baseline estimation, albeit approximate. Although modeling noise can affect feedback reliability, the rigid-body transformation adheres to fundamental motion principles and generalizes across robotic systems. While this model simplifies nonlinear complexities, the RL part in our framework compensates effectively within an informative closed-loop system.
IV-D3 Model simplification
TABLE V: Evaluation of the internal predictive model’s complexity, tested on quadrotors tracking zigzag trajectories.
Model Train Eval
static disturbed static disturbed
Simplified 0.057 0.064 (+12.28%) 0.101 0.119 (+17.82%)
Full 0.061 0.067 (+9.84%) 0.095 0.105 (+10.53%)
We investigate the impact of predictive model simplification on performance. A quadrotor serves as our testbed due to its well-studied dynamics model [40]. We compare our approach with a variant that uses the full-dynamics model for tracking zigzag trajectories. As shown in Tab. V, the results indicate that despite the simplification, performance degradation is marginal. Notably, this simplification improves generalization across different morphologies, making it applicable to complex systems that are difficult to model.
@@ V Conclusion
In this paper, we propose Neural Internal Model Control (NeuralIMC), a framework that bridges model-based control with RL-based neural control. By leveraging body dynamics and a white-box rigid-body transformation, we avoid high-dimensional nonlinearity, using a simplified predictive model in a closed-loop feedback structure. Model-free RL enables the policy to react to predictive errors. NeuralIMC outperforms state-of-the-art controllers on quadrotors and quadrupedal robots. Real-world tests on a heavily loaded quadrotor demonstrate robustness to disturbances and dynamic variations, highlighting its potential for interpretable and adaptive learning-based control. For the future work, we’d like to conduct more expanded real-world experiments includes testing on diverse robots and improving disturbance training [45].